3. (30 points) Critically-damped - Determine vo(t)! 200 10uH Vo(t) i(t) v(t) 1Vu(t) 100nF
Answer #1

zon lo KH ICS) = _ils L R+ SL + Śc + 1 loon = UCL), sze + SRC+1 26 SL VL DRI() + 16 Fść š + ßst the 1 s² +6x166) s + 1012 letVGS) = T(S). 102 - S(SH1082 - to forta VCS) = - Stock - 106 X = (1 - 2107_606e-loft) ecke), y (SH)2